OpenStudy (anonymous):

Solve the system of equations using any method X=4y 3x-2y=20

OpenStudy (australopithecus):

well we know x = 4y so, 3(4y) - 2y = 20 solve for y then sub the value of y into the first equation to solve for x

OpenStudy (anonymous):

12-2y=20

OpenStudy (australopithecus):

where did the y go?

OpenStudy (anonymous):

12y-2y=20

OpenStudy (australopithecus):

do you not know how to solve from that point for y?

OpenStudy (australopithecus):

Just factor out the y

OpenStudy (anonymous):

10y=20 y=2?? is this corect

OpenStudy (australopithecus):

yup

OpenStudy (australopithecus):

now sub it into one of the original equations and solve for x and you have your solution

OpenStudy (anonymous):

y=2 x= 8

OpenStudy (australopithecus):

the point where the two lines intersect :)

OpenStudy (anonymous):

is this corect??

OpenStudy (australopithecus):

yes it is correct